Two cars leave towns 400 kilometers apart at the same time and travel toward each other. One car's rate is 18 kilometers per hour less than the other's. If they meet in 2 hours, what is the rate of the slower car?
Do not do any rounding.

Answer by mananth(16946)

one ar speed = x
other car speed = x-18
..
they are travelling towards each other.
so combined speed = x+x-18= 2x-18
..
t=d/s
2= 400/(2x-18)
divide by 2
1= 200/(2x-18)
2x-18=200
2x=200+18
2x=218
/2
x=109 speed of one car
speed of other car === > x-18===> 109-18=91mph